This year, Bosch Global will produce 60,000 commercial ovens at its plant in New Jersey. There is a concern that this output may not meet demand. To produce the 60,000 commercial ovens, each production worker at the plant will work 180 hours per month. If the labor productivity at the plant is 0.12 ovens per labor-hour, how many production workers are employed at the plant?

Masters level class- need work with Answers
Answer #1

Yearly output required = 60000 ovens

So, monthly output required = 60000/12 = 5000 ovens

No. of ovens produced per labor-hour = .12

So, no. of labor hours required to produce 5000 ovens = 5000/.12 = 41666.67 labor hours

Number of hours per worker per month = 180

So, number of labor per month required = 41666.67/180 = 231.48 or 232 workers

Therefore, 231.48 or 232 workers will be required per month to complete the task of producing 60000 overs in a year.